OpenStudy (anonymous):

Find the values of x+y if x^{2} + y^{2} =36 and xy=-10.

jimthompson5910 (jim_thompson5910):

Hint: (x+y)^2 = x^2 + 2xy + y^2

OpenStudy (anonymous):

so how do you go about solving the problem?

jimthompson5910 (jim_thompson5910):

(x+y)^2 = x^2 + 2xy + y^2 (x+y)^2 = (x^2 + y^2) + 2(xy) How about that?

OpenStudy (anonymous):

ohh okay so it ends up being 16=(x+y)^2 and then simplifying it becomes 4=x+y thanks

jimthompson5910 (jim_thompson5910):

very good